Operation Summary for December 20, 2010 - December 22, 2010 |
Monday Dec. 20 The day shift began with Operations monitoring the Tevatron (TEV) store 8377 (271.01E30) and the Recycler stash (51.7E10), with the Antiproton Source (Pbar) stacking (21.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-994 (JASMIN) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported that regulation problems with two Linac quadrupole magnet power supplies (L:Q51 and L:Q53) resulted in downtime between 10:47 AM and 12:27 PM. During the evening shift, Operations disabled the T-944 beam at 6:15 PM. Operations terminated store 8377 at 10:52 PM and put the TEV into HEP shot setup. Tuesday Dec. 21 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (5.3E10), with Pbar stacking (42.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 8378 at 12:04 AM with an average initial luminosity (AIL) of 361.5E30. Operations established beam to T-944 at 4:23 AM. During the day shift, Operations terminated store 8378 at 8:33 AM to allow accesses to TEV sectors A-E and the collision halls while cryo system personnel worked on the TEV sector D1 wet engine. Operations reported that the TEV accesses concluded, but the TEV quenched at 3:14 PM during its checkout. During the evening shift, operators had the TEV on and in a dry squeeze by 5:46 PM. Operations disabled the T-944 beam at 6:05 PM; this ended JASMIN’s run. Operations conducted TEV orbit checks. Operations put the TEV into shot setup and established store 8383 at 11:26 PM with an AIL of 367.1E30. Wednesday Dec. 22 The midnight shift began with Operations monitoring store 8383 (287.94E30) and the stash (153.4E10), with Pbar stacking (19E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations monitored the accelerator complex. The Plans for Wednesday The plans for today are to stack, stash, and store. However, the Run Coordinator will terminate store 8383 this morning to allow D0 to access and work on their forward muon detector. He will also allow RF personnel to work on a Booster RF station, cryo system technicians to work on the Pbar wet engine, NuMI personnel to conduct target scans, Water personnel to bring Column #2 on line at CUB, and for MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 12/22/2010 email Fermilab |
| Security, Privacy, Legal | |